Circular No. 8682

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 CK_{10} (CATALINA) C. W. Hergenrother, Lunar and Planetary Laboratory, reports that his

CCD frames taken (with A. C. Smith) at the Catalina 1.54-m Kuiper telescope on Feb. 24 and 25 UT, and at the Vatican 1.8-m Lennon telescope on Feb. 28.5, show that 2006 CK_{10} — reported as asteroidal when discovered by the Catalina Sky Survey (cf. MPEC 2006-C44; discovery observation tabulated below) — has a cometary appearance; a co-added 1440-s R-band exposure on Feb. 24.48 shows a condensed coma 14′′ in diameter and a broad tail 20′′ long in p.a. 210^{o}. E. J. Christensen reports that four Catalina Sky Survey co-added 90-s CCD images obtained by A. R. Gibbs and himself with the 0.68-m Schmidt telescope on Mar. 2.35–2.36 show a faint coma ≈ 20′′ in diameter (mag 17.3), elongated toward p.a. 200^{o}.

2006 UT  α_{2000}  δ_{2000}  Mag.
Feb. 4.43262  14^{h}16^{m}00.^{s}15  +35^{o}00′10′′.5  19.0

Additional astrometry, the following parabolic orbital elements, and an ephemeris appear on MPEC 2006-E04.

T = 2006 July 2.6713 TT  ω = 143.^{o}1050
Ω = 243.7873 }2000.0
q = 1.757046 AU  i = 144.2441

RS OPHIUCHI A. Evans, Keele University; T. Kerr, Joint Astronomy Centre; Y. Mat-

suoka and Y. Tsuzuki, University of Tokyo; T. R. Geballe, Gemini Observatory; R. D. Gehrz and C. E. Woodward, University of Minnesota; M. F. Bode, Liverpool John Moores University; T. J. O’Brien and R. J. Davis, University of Manchester; J. P. Osborne and K. L. Page, University of Leicester; G. Schwarz, West Chester University; S. Starrfield and J.-U. Ness, Arizona State University; J. Krautter, University of Heidelberg; J. Drake, Smithsonian Astrophysical Observatory; and S. P. S. Eyres, University of Central Lancashire, report that a spectrum (range 880 nm–2.5 μm) was obtained with the 3.8-m U.K. Infrared Telescope (+ UIST) on Feb. 24.64 UT. The spectra show strong Paschen lines from 4-3 to 12-3, Brackett lines from 7-4 to 20-4, Pfund lines from 17-5 to 23-5, He I at 1.083, 2.058, and 2.112 μm, O I 3P-3D^{0} at 1.128 μm (cf. IAUC 8673), and other lines that are tentatively identified with O I, O II and C I. The FWHM of the resolved Paα line is 1130 ± 50 km/s, but with wings extending to 3000 km/s.

2006 March 2  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green